OpenStudy (anonymous):

ONLY @brandon10995 I AM TRYING TO HELP HIM SO IF YOU ARE NOT BRANDON, I LOVE YOU BUT BYE!!

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

3x-6y+2=0 was your question right?? @Brandon

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I will walk you through! so you have something equal to 0. you can always move the numbers over to the other side of the equality sign. So in this case, you could do the followin; 3x-6y+2=0 subtract 2 from each side. 3x-6=-2 does this seem familiar now? and now add 6 to each side. 3x=4. now divide 3 each side. x=4/3

OpenStudy (anonymous):

your welcome brsndon. xD

OpenStudy (anonymous):

so you basically act like the 0 isn't there? I need to find the slope and y intercept of the line

OpenStudy (anonymous):

Isent the slope 4/3?

OpenStudy (anonymous):

if i wasn't stuck i wouldn't be asking right now haha